### v1

Abstract
Large language models (LLMs) are revolutionizing every aspect of society. They are increasingly used in problem-solving tasks to substitute human assessment and reasoning. LLMs are trained on what humans write and thus prone to learn human biases. One of the most widespread human biases is overconfidence. We examine whether LLMs inherit this bias. We automatically construct reasoning problems with known ground truths, and prompt LLMs to assess the confidence in their answers, closely following similar protocols in human experiments. We find that all five LLMs we study are overconfident: they overestimate the probability that their answer is correct between 20% and 60%. Humans have accuracy similar to the more advanced LLMs, but far lower overconfidence. Although humans and LLMs are similarly biased in questions which they are certain they answered correctly, a key difference emerges between them: LLM bias increases sharply relative to humans if they become less sure that their answers are correct. We also show that LLM input has ambiguous effects on human decision making: LLM input leads to an increase in the accuracy, but it more than doubles the extent of overconfidence in the answers.
Large language models (LLMs) are revolutionizing every aspect of society. Among other things, they are used in problem-solving tasks that require careful reasoning and assessment. While LLMs are well-known to reliably reproduce knowledge on which they have been trained (Zhao et al., 2023), they are prone to making mistakes in reasoning tasks on which they have not been directly trained (Li et al., 2024b; Huang and Chang, 2023; Chang et al., 2024; Xu et al., 2023; Pawitan and Holmes, 2025).
LLMs are trained on what humans write. As such, they are prone to develop human biases in their reasoning. One of the most prevalent biases in human judgment is overconfidence (Hoffrage, 2022; Kahneman, 2011; Moore and Healy, 2008; Malmendier and Taylor, 2015). Humans are prone to underestimating the limits of their knowledge (Dunning, 2011; Kruger and Dunning, 1999), leading them to overestimate the accuracy of their knowledge. Evaluating the correctness of their reasoning in different domains may also lead to the conjunction fallacy (Tversky and Kahneman, 1983), amplifying the overall overconfidence in judgment.  [^1] Various non-cognitive factors have also been shown to contribute to overconfidence, such as social signaling, (Burks et al., 2013), environmental factors (Goette et al., 2015), or strategic considerations to feign confidence (Charness et al., 2018) that may also provide an evolutionary mechanisms for the trait (Johnson and Fowler, 2011; Marshall et al., 2013).
Overconfidence has been shown to affect decision in many areas. In the professional domain, overconfident managers are more likely to use equity financing (Malmendier and Tate, 2005; Malmendier et al., 2011) and more likely to engage in unprofitable merger (Malmendier and Tate, 2008), exhibiting little awareness or learning over time (Huffman et al., 2022). There is also evidence of overconfidence affecting everyday behaviors, such as working out regularly or dietary choices (Arni et al., 2021), or financial investments (Barber and Odean, 2001).
In this context, LLMs may hold, both, promise and peril. On the one hand, overconfident individuals may be challenged in their reasoning by LLM input, thus mitigating human bias. On the other hand, if overconfidence manages to enter LLMs, this may exacerbate confidence in human decision makers.
The accuracy and overconfidence of LLMs
In this paper, we examine whether LLMs inherit overconfidence when evaluating their own reasoning, and whether this affects human decision makers. While conventional wisdom starts from the "bias in, bias out" hypothesis (Barocas and Selbst, 2016; Mayson, 2019), this need not always hold: theoreitcal work shows that overconfident agents learning from mis-specified models have complex and sometimes counterintutive implications heidhues2019overconfidence. It is possible that biased training data leads AI to be less biased towards that group (rambachan2020bias). 222rambachan2020bias show that in many realistic environments, bias in training data, stemming, e.g. from human bias against demographic group, can lead to less biased behavior of an AI towards that group. However, a corollary of their result is that biased training data will still affect other groups, with the AI being more biased towards them, even though humans don’t display bias towards those groups. To our knowledge, no formal results exist for the specific context of AI or LLM reasoning when trained on biased data. We thus test for overconfidence in LLM reasoning, and probe into the underlying mechanisms. We ask whether the bias is amplified or weakened relative to humans performing the same task. We prompt five commonly used LLMs with a large number of reasoning problems to which we know the true answer. The questions are generated using the algorithm in li2024drowzee to generate questions on which the LLMs are unlikely to be trained on. Thus, to arrive at an answer, the LLMs need to engage in reasoning on the spot.  [^3] Several recent papers have examined measures of calibration that can be extracted from LLMs’ intermediate outputs. Prior work has taken three main approaches to measuring LLM confidence: (1) logit-based estimation, which requires internal model access (yin2023large; gpt4report2023; he2023investigating; zhang2024calibrating); (2) direct confidence elicitation through prompting (tian2023just; xiong2024can; wei2024measuring; wen2024mitigating); and (3) auxiliary model approaches, ranging from single-model prediction (ulmer2024calibrating)to multi-source integration methods (zhao2024pareto). These calibration studies predominantly rely on standard question-answering datasets, and it is likely that the LLMs have been trained on a large fraction of the questions used (tian2023just; wei2024measuring; zhang2024calibrating; xiong2024can). Thus, these studies do not address LLMs’ confidence in questions that require them to reason. By contrast, we utilize algorithmically generated questions that guarantee minimal training contamination, allowing us to study confidence in answers that involve reasoning by LLMs. We ask the LLMs to assess the confidence in their answers, closely following similar protocols in human experiments. We elicit the LLM’s confidence in the correctness of the answer, the facts used, and the reasoning.
We find that all five LLMs we examine are overconfident (Table 1). The LLMs on average overestimate the probability that their answer is correct between 20% (for GPT o1) and 60% (for GPT 3.5), with the other models falling somewhere in between. We find that more advanced models such as GPT 4o or GPT o1 have higher accuracy than GPT 3.5 or Llama 3.2, in line with earlier findings (shahriar2024putting; zhong2024evaluation; dubey2024llama). Interestingly, the confidence judgments across models are very similar, and in no way reflect the large differences in accuracy rates.
Next, we examine how, within each LLM model, accuracy is related to confidence across the different questions. Panels A and B of Figure 1 show the accuracy rates as a function of confidence, for the GPT and Llama models, respectively. The panel shows a strong and positive association between accuracy and confidence. The graphs also show that even when a model is fully confident in its answer, there is still substantial bias: Panel A of Table 2 shows that the bias varies between 14% and 20% for the top performing models (GPT 4o, o1, and Llama 3.1), and even more for GPT 3.5 and Llama 3.2.
For all models except Llama 3.2, bias increases with lower confidence levels. The estimates in Table 2 quantify this effect: a 10% drop in confidence is associated with a drop in accuracy of around 25% for GPT 4o and GPT o1, and around 15% for Llama 3.1 (see panel A in Table 2). This implies that the bias gets larger as the models’ confidence declines.
The strong association between confidence and accuracy also extends each component of the reasoning process in which the LLMs have to engage. For each question, we separately elicit the the LLM’s confidence that the facts used are correct, and that the reasoning applied is correct. These confidence measures are highly correlated with confidence that the overall answer is correct (correlations range from 0.49 to 0.92 across models , see Table D.7 and Table D.7 in the SI). We also calculate the semantic similarity between the facts used and the reasoning applied with the sufficient set of facts, and confidence that the reasoning is correct (see section A.3 in the SI for details). We find significant associations between confidence and similarity for, both, facts and reasoning ($p<0.001$ in all cases, see Tables D.7 and D.7 in the SI). Thus, there is a palpable, but biased, relationship between accuracy and confidence in the answers, and this association extends to the necessary components for the reasoning.
We also examined the stability of these results. In five replications of the results for GPT 3.5 and 4o, we retrieve nearly identical results with regard to the answer and the confidence. The answers change in less than 2 percent in each iteration. The changes we observe also do not have a clear direction: overall, accuracy only increase by 1.7 percent after five replications (see section D.8 in the SI).
LLM vs. human overconfidence
We compare the accuracy, confidence, and Dunning-Kruger effect of LLMs to a human benchmark. In a large-scale online experiment on Prolific, we recruit participants to answer 10 questions of a subset of 2000 (see SI section B). Like the LLMs, the human participants have to pick one of the answers (Yes or No), and indicate the probability with which they believe that their answer is correct. Human participants answer 66% of the questions correctly. This puts them on par in terms of accuracy with GPT 4o or Llama 3.1, but slightly behind GPT o1. Their confidence is significantly lower than that of any of the LLMs: human participants think their accuracy is 70%, thus overestimating it by only 4% on average ($p<0.01$, see Table S2 in the SI). This is far lower overconfidence than any of the LLMs with comparable accuracy. There is a maked difference in shape of the distribution of confidence (see Figure S5 in the SI): compared to LLMs, humans’ confidence appears far more nuanced, and with a mode (nearly 35% of the answers) stating that it is 50:50 whether their answers are correct.
Panel C of Figure 1 shows the relationship between accuracy and confidence in the human sample, compared to the three most accurate LLMs. Two striking features emerge from the graph: first, when humans are 100% sure that their answer is correct, their accuracy rate is significantly lower at 81% ($p<0.01$, see column 1 in Table S11 ). Thus, for questions that humans consider relatively easy and are thus 100% sure of their answer, they are highly overconfident. This level of overconfidence is roughly on par with the same conditional accuracy rates of GPT 4o or Llama 3.1. However, while originating from the same level of bias when 100% sure of their answers, the confidence gradients between LLMs and humans are sharply different. As discussed before, LLMs become more biased as they enter more uncertain territory. In sharp contrast, the human confidence gradient is only about 0.5, indicating that human bias decreases as humans become less certain. As can be seen in the figure, humans eventually end up slightly underconfident: when participants think their accuracy is only 50%, it is actually slightly higher (54%, significantly higher than they expect $p<0.01$).
We interpret this sharp difference in the confidence gradient as a stark manifestation of the Dunning-Kruger effect (dunning2011dunning; kruger1999unskilled). Our test cases require LLMs to reason to answer them, as they are unlikely to be trained on the answer. This may cause LLMs being "trapped" in their prediction model in a way that is hard for them to overcome: the model-internal mechanisms cannot provide them with a sense of what knowledge is not part of their training data. The LLMs pick the answer that is most consistent with what they were trained on, and form an estimate of their accuracy based on what is in their training data. As LLMs move into less certain territory, this problem gets stronger, and causes LLMs to have much less of a sense of the limits of their knowledge. Human participants may recognize in a question, e.g., that one of the names is a Roman emperor, and realize that she does not know much about Roman history. By contrast, in an LLM, there is no mechanism to tell it that there might be other Roman public figures that it wasn’t trained on. This results in overconfidence far stronger than what is observed in humans though the mechanisms of the Dunning-Kruger effect. Interestingly, while more advanced models like GPT 4o and GPT o1 have higher accuracy rates for test cases where they are certain their answer is correct, they also display a significantly stronger confidence gradient, and, hence a stronger Dunning-Kruger effect.
Exposure to LLM input
While the computer science literature has documented limitations in LLMs’ ability to reason, human users may not be aware what type of questions LLMs have been directly trained on, and which questions involve reasoning by LLMs. In a second experiment, we test how human performance, in terms of, both, accuracy and bias, is affected by exposure to LLM answers and LLMs’ confidence in their answers. We test this in an information-intervention design: like in the previous experiment, participants answer questions and state their estimate of the probability of their answer being correct. In a second stage, participants in the "LLM Answer" condition are then provided with the answer that the LLM picked for this question. We randomly choose to show the answer from, either GPT 4o, o1, or Llama 3.1, i.e. only the three models with the highest accuracy. Participants are then given the opportunity to revise their own answer and confidence in the answer. In the "LLM Answer + Confidence" condition, participants are shown the LLM’s answer as well as the LLM’s confidence in its answer. In a control condition, participants are simply given the opportunity to revise their answer (see SI section C for more details).
Figure 2 shows how exposure to LLM input changes the accuracy (Panel A) and bias with regard to accuracy (Panel B). Exposure to, either, LLM answer or answer plus confidence increases the accuracy of the human participants’ answers between 5.6 and 7 percentage points. However, both conditions the participants’ confidence in their answers by even more, increasing bias by 4.2 percentage points in the answer condition, and 7.6 percentage points in the Answer + Confidence condition. Displaying the LLM’s confidence does not lead to a significant increase in accuracy, but increases bias significantly more than only providing the LLM answer ($p<0.01$, see SI Table S12). The amount of bias that exposure injects into humans’ assessments is large: at baseline, participants exhibited only moderate overconfidence, of around 4%. Thus, exposure to LLM input doubled (in LLM Answer), or nearly tripled (in LMM Answer + Confidence) their bias.
We next examine which types of users of LLMs are driving these overall treatment effects. LLMs help improve the accuracy of individuals with low baseline confidence in their answer, raising it by 8.6 and 11.9 percentage points, respectively (see Panel A of Table 3. However, LLM exposure also substantially increases bias: human confidence in the revised answer increases by far more than what is justified by the increase in accuracy. Bias increases by 7.0 percentage points in the LLM Answer condition, and by 14.1 percentage points when, in addition, the LLM confidence is also displayed. The LLM answer nearly triples bias, and displaying the LLM confidence more than quadruples it in questions where individuals have below-median confidence at baseline. Interestingly, displaying LLM confidence is not necessary to increase bias in humans. By contrast, individuals with above-median confidence in their answer experience seem ‘immovable:’ they experience neither a gain in accuracy nor a increase or decrease in bias.
We also explore how heterogeneity in LLMs’ confidence affect exposure to LLM input. Table 3 shows the results. Panel B shows how low-confidence (80% to 90%) and high-confidence (90% - 100%) answers from LLMs affect the users accuracy and bias. Low-confidence answers from the LLM do not increase accuracy (the point estimate is slightly negative) in either condition. However, bias substantially increases. By contrast, answers in which LLMs have more than 90% confidence increase accuracy, but fail to decrease bias.
Discussion
Overall, our results suggest that LLMs are far more overconfident than humans in tasks that involve LLM reasoning. Humans, in turn, do not sufficiently discount LLM advice. On average, exposure to LLMs input leads humans to become more accurate, but also more biased. The analysis in sub-sambples shows that this effect is driven by humans with low baseline confidence in their answer for, both, the accuracy and bias.
While LLMs generate output that is highly valuable to users by tapping into answers to questions on which they are pre-trained, our analysis shows tasks that require LLMs to reason may not increase human welfare. We examine this formally in section E of the SI. We set up a model in which payoffs depend on, both, the probability of getting the answer correct, and a costly investment that an individual makes. If the payoffs are largely proportional to the probability of giving a correct answer, then LLMs increase welfare. In such a setting, bias per se does not have any costs. However, many tasks also require a decision or investment depending on the assessed probability of success. In these environments, overconfidence leads to over-investment, and reduces expected payoffs relative to what would be optimal. It is easily possible that the increase in bias offsets the gain in accuracy (see Result 3 in SI section E). Quite intuitively, the more elastic investment responds to the perceived probability of success, the greater the scope for increased bias from LLM exposure to reduce welfare. The model also makes more subtle predictions. Suppose, as our data suggest, that individuals are biased at baseline. In this case, LLM exposure can lead to lower welfare even if only the probability of providing a correct answer increases. The reason is that the higher accuracy increases investment. Because marginal costs of investment are increasing, this makes the over-investment from bias more costly. As we show, this effect can be so strong that it offsets the gains from higher accuracy (see Result 2 in SI section E). This highlights an important warning: even intuitive indicators such as increased accuracy from LLM exposure may be not be indicative of higher payoffs. The opposite may, in fact, be the case, as our examples demonstrate.
In assessing the benefits from LLMs, is also noteworthy that LLMs and human participants express uncertainty about different questions. In our sample, the correlation between average human confidence and average LLM confidence is only 0.082, see SI Table S6. This tends to improve welfare, as individuals who are uncertain about an answer receive roughly average-quality LLM answers, which have a higher accuracy than humans who are uncertain at baseline. If uncertainty were more correlated, LLM input would lead to lower accuracy for low human baseline confidence, but bias would increase substantially (see SI, Table S7.) Correspondingly, high-confidence LLM answers would be ‘wasted’ on high-confidence humans, where little improvement in accuracy is possible thus confirming their answers and further increase their bias.
The implications of LLM overconfidence span practical, ethical and technical dimensions across stakeholder groups. Our results provide users with a new benchmark for how users of LLMs should approach reasoning from LLMs, counteracting the widespread overestimation of LLM capabilities (KLINGBEIL2024108352; holbrook2024; info:doi/10.2196/56764). Secondly, our results provide guidance for the future development of LLMs. Architectural choices in training objectives currently prioritize fluency over accuracy, incentivizing fabricated but coherent responses (yin2023large; fi15100336). The lack of a built-in uncertainty correction mechanism generates hallucinated responses and overconfidence (shorinwa2024survey). Current evaluation metrics (wei2024measuring; geng-etal-2024-survey) fail to capture calibration quality, necessitating new frameworks for quantifying and gauging overconfidence such as the one we presented here. Developing novel interpretability frameworks (wen2024from; 9857619) that pinpoint how and why overconfident and biased predictions emerge can support clearer communication of model uncertainty and limitations to users.
Our results indicate that making LLMs more knowledgeable (in terms of training data, or parameters in the model) is unlikely to remove overconfidence: while we find that overconfidence decreases with knowledge when LLMs are certain to be correct, they also show that the Dunning-Kruger effect is significantly stronger for larger LLMs as uncertainty increases. In other research, we examine the prevalence of framing effects in LLM answers fu2025llms. We find strong framing effects in all models. Larger and newer models display weaker framing effects, but the gains appear to level off: for instance, we find no significant reduction in framing effects between OpenAI’s o1 and and the most recently released o3. This points to the importance of incorporating validation mechanisms to check the LLM’s reasoning process (see, e.g. binder2024looking), rather than simply trying to make models larger. This line of research has the potential to stimulate new algorithmic approaches and training strategies, ultimately fostering accountable, trustworthy, and ethically sound AI systems (NEURIPS2022_ff887781; li2024overconfident).
More broadly, this research highlights the usefulness and importance of evaluating the presence, extent and consequences of behavioral biases in LLM reasoning. There are many other potential biases well-known in behavioral science, such as basic violations of rationality (chen2023emergence), framing effects (tversky1981framing; kahneman1984choices), gender bias (sun2019mitigating), the conjunction fallacy (tversky1983extensional; busemeyer2011quantum), or conformation bias (klayman1995varieties; nickerson1998confirmation) that permeate training data generated by humans. Behavioral science has established paradigms to measure these departures from rationality. A nascent literature examines how individuals respond to (potentially biased) AI input. Fruitful collaboration with computer science awaits to explore the overarching mechanisms of how they manifest in LLMs, and how they can be reduced.
Tables and figures
Model a) Fraction of b) LLM confidence c) Bias calculated
correct answers in answer ($\widetilde{p}$) based on $\widetilde{p}$
gpt 3.5 0.35 0.94 0.59
(0.045) (0.473)
gpt 4o 0.63 0.94 0.30
(0.053) (0.470)
gpt o1 0.73 0.95 0.22
(0.044) (0.433)
Llama 3.1 0.63 0.86 0.23
8B (0.102) (0.466)
Llama 3.2 0.61 0.94 0.33
3B (0.075) (0.486)
Table 1: Accuracy, confidence, and bias across models
  • Notes: $N=10,000$. Standard deviations in parentheses (divide by 100 to obtain approximate SE of means). All accuracy rates and confidence measures are significantly different between models. All biases are significantly different from zero. ($p<0.001$, see SI Table S9.)
gpt 3.5 gpt 4o gpt o1 Llama 3.1 Llama 3.2
Confidence 1.36 2.57 2.77 1.46 0.67
gradient (0.10) (0.09) (0.13) (0.05) (0.07)
Predicted accuracy 0.43 0.80 0.86 0.83 0.65
if $\widetilde{p}=1$ (0.01) (0.01) (0.01) (0.01) (0.01)
$N$ 9,795 9,973 9,990 9,750 9,910
$R^{2}$ 0.01 0.07 0.07 0.08 0.01
Table 2: The confidence gradient in accuracy
  • Notes: Dependent variable is correct answer (= 1). The table presents the gradient of accuracy with respect to self-reported confidence in answer correctness across models. Confidence levels are normalized by subtracting 1 from the original confidence scores. Regression is weighted by the frequency of confidence levels. All coefficients are significant at p<0.01.
Human Baseline Confidence LLM Confidence
$\Delta$ Accuracy $\Delta$ Bias $\Delta$ Accuracy $\Delta$ Bias
LLM Answer 0.086 0.070 -0.033 0.133
(0.015) (0.017) (0.018) (0.020)
LLM Answer + Conf. 0.119 0.141 -0.028 0.160
(0.017) (0.018) (0.020) (0.021)
LLM Answer -0.064 -0.063 0.136 -0.138
$\times$ High Confidence (0.018) (0.019) (0.020) (0.021)
LLM Answer + Conf. -0.091 -0.124 0.153 -0.119
$\times$ High Confidence (0.019) (0.019) (0.024) (0.024)
Constant 0.003 0.018 -0.002 0.014
(0.004) (0.005) (0.004) (0.005)
Observations 11,610 11,610 10,957 10,957
R2 0.01082 0.01700 0.02477 0.01915
Table 3: Subsample Analysis of LLM Exposure Experiment
  • Notes: LLM confidence is mean-centered at 0.8; Baseline human confidence is mean-centered at 0.5; High Confidence is defined as confidence above 0.9 for LLM and above the median for human baseline. The regressions also control for high confidence indicators, though these results are not displayed in the table. Standard errors in parentheses are clustered at the participant and question levels. Illustrations can be seen in SI Figure S4.
Figure 1: LLM Confidence and Accuracy
Notes: Panel A shows the relationship between accuracy and confidence for GPT models, and Panel B shows the same relationship for Llama models. Panel C compares the accuracy-confidence relationship between advanced LLM models and human responses on the question subset for the human baseline benchmark. The size of circles in each panel scales with the number of observations, and all slopes are significant ($p<0.001$ in all cases; see Table 2 for Panel A and B, and Table S11 in Supplementary Information for Panel C). Shaded areas indicate 95% confidence bands.
Figure 2: Treatment Effects on Changes in Accuracy and Bias
Notes: Panel A shows the treatment effects on change in accuracy across experimental conditions. Panel B shows the treatment effects on the change in bias across the same conditions. Error bars represent 95% confidence intervals. For detailed regression analysis, see SI Table S12.
Appendix A The LLM Experiment
A.1 Question Generation
A.1.1 Triple Extraction from Wikipedia
The first step involves extracting fundamental facts from the input knowledge data into structured triples for logical reasoning. Wikidata is a structured knowledge base that provides a vast collection of interconnected data on Wikipedia, including entities and their relationships, making it an ideal resource for extracting structured triples. We select the top-ten popular categories from Wikidata’ entity and relation classifications to guide the extraction of facts (for category details, see Table S14). Each fact is formatted as a triple (subject, predicate, object), where the subject and object represent entities, and the predicate describes their relationship. The extraction is done by searching through the knowledge base, retrieving all facts related to each entity and its corresponding relations. This methodical organization of triples establishes a solid foundation for subsequent logical reasoning operations.
A.1.2 Logical Reasoning Process
Step 1: Rule Generation.
In this step, an automatic rule generator is designed to iterate its predicates and generate the derivation rules $Q$ according to the relation category. There are five reasoning types of our interest: Negation Reasoning, Symmetric Reasoning, Inverse Reasoning, Transitive Reasoning, and Composite Reasoning. The definition and organization of facts using each reasoning type is detailed in the following paragraphs.
Rule#1: Negation Reasoning. Based on a given factual knowledge, we can determine whether the opposite of this fact is correct or incorrect by applying Definition 1.
Definition 1.
Negation Reasoning Rule $[Neg]$. Given a factual knowledge triple $(s,\mathit{nm},o)$, then we can derive the new triple $(s,\overline{\mathit{nm}},o)$ is not valid. $\overline{\mathit{nm}}$ indicates the negation of the relation $\mathit{nm}$.
[…]
With this rule, from the triple (Haruki Murakami, did not win, the Nobel Prize in Literature in 2016), we derive that the negation of this triple (Haruki Murakami, won, the Nobel Prize in Literature in 2016) contains false factual knowledge.
Rule#2: Symmetric Reasoning. In symmetric relations, if the subject and object in a triple maintain coherence upon interchange, a new triple can be deduced in accordance with Definition 2.
Definition 2.
Symmetric Reasoning Rule $[Sym]$. Given a factual knowledge triple $(s,\mathit{nm},o)$, then we can derive a new triple $(o,\mathit{nm},s)$.
[…]
With this rule, from the original triple (Haruki Murakami, different_from, Haruki Uemura), we derive a new triple (Haruki Uemura, different_from, Haruki Murakami) (Haruki Uemura is a Japanese judoka). Note that the symmetric reasoning rule is primarily utilized within the composition reasoning rule (to be detailed next) and does not introduce new knowledge on its own.
Rule#3: Inverse Reasoning. In an inverse relation, the subject and object can be reversely linked through a variant of the original relation, as defined in Definition 3.
Definition 3.
Inverse Reasoning Rule $[Inverse]$. Given a factual knowledge triple $(s,\mathit{nm},o)$ and a reversed relation $\mathit{nm}^{\prime}$ of $R$, then we can derive a new triple $(o,\mathit{nm}^{\prime},s)$.
[…]
Rule#4: Transitive Reasoning. In transitive relations, if the object in one triple is the subject of the second triple, we can therefore derive a new triple following the Definition 4.
Definition 4.
Transitive Reasoning Rules $[Trans]$. Given two factual knowledge triples $(s_{1},\mathit{nm},o_{1})$ and $(s_{2},\mathit{nm},o_{2})$, if $o_{1}$ is semantically equivalent to $s_{2}$, then we can derive a new triple $(s_{1},\mathit{nm},o_{2})$.
[…]
An example here is:
$\frac{\begin{array}[]{c}loc\_in(s_{1},o_{1}),\leavevmode\nobreak\ loc\_in(s_{2},o_{2}),\leavevmode\nobreak\ o_{1}=s_{2}\\ \end{array}}{\begin{array}[]{c}loc\_in(s_{1},o_{2})\end{array}}[Trans].$
With this rule, from triples (Haruki Murakami, locate_in, Kyoto) and (Kyoto, locate_in, Japan), we derive a new triple (Haruki Murakami, locate_in, Japan).
Rule#5: Composite Reasoning. The previous four reasoning rules are all meta-rules capturing the most basic and fundamental logical relations among the facts and rules. Several basic reasoning rules can be chained together to form a composition reasoning rule if the relations in the rules are logically related. Composite reasoning rules can generate knowledge that requires multiple steps of reasoning.
Definition 5.
Composite Reasoning Rules $[Comp]$. Given multiple basic reasoning rules or predicates $[Rule_{i}]\in\{[Neg],[Sym],[Inverse],[Trans],[Predicates]\}$, we can chain them up to form a new composite reasoning rule.
[…]
Step 2: Prolog Inference.
With the predetermined rules, we can be assisted with the Prolog engine, asserting all the related triples and consulting the reasoning rules. We use $\llbracket R\rrbracket_{\mathcal{P}}$ to denote the query results of $R$ w.r.t the Prolog program $\mathcal{P}$. When $R$ contains no variables, it returns Boolean results indicating the presence of the fact; otherwise, it outputs all the possible instantiations of the variables. Then by obtaining solutions from Prolog queries, we can generate new knowledge triples based on the entities and their relations provided. For each instantiation that contains one subject “s” and one object “o”, we then compose them with the new predicate to form the newly derived triple. They are later used as the complex knowledge to generate test cases and the corresponding oracles.
A.1.3 Predicate Verification Process
From the new knowledge triples, we manually review the predicate component to ensure the accuracy and validity later in the generated question-answer pairs. After removing potentially ambiguous knowledge, we ultimately retain 476 predicates along with their corresponding triples.
A.1.4 Question Formulation and Refinement
After generating triples from the Wikipedia database and manually verifying predicate verbs, we utilize GPT-4o to automatically create corresponding initial questions. To enhance the naturalness of these questions and minimize potential misinterpretations by generative AI in subsequent steps, we employed a sophisticated prompt for question refinement. This prompt was designed to check for grammatical errors and improve overall question formulation. The full prompt is as follows:
Please formulate a question of the form “Is it true that …”, using the following structure: the subject is “[subject]”, the object is “[object_]”, and the predicate in short-hand notation is “[predicate]”. Before generating the final question, please perform the following checks and adjustments:
  1. 1.
    Predicate Validity Check: Examine whether the given predicate “[predicate]” makes grammatical sense in English. If it doesn’t, please reformulate it to ensure it follows proper English syntax and conventions. For example, same_instance_of might be reformulated to “is the same instance as” or “is identical to”.
  2. 2.
    Sentence Completeness Check: Assess whether the current predicate is sufficient to generate a complete, coherent sentence. If not, extend it appropriately without altering its core meaning. For example, same_named_after could be extended to “is named after the same person/thing as”.
After making these adjustments, generate a grammatically correct and complete question that accurately represents the relationship between the subject and object using the (potentially modified) predicate. Your final output should be a single, well-formed question in the format “Is it true that [subject] [predicate] [object]?”, where [predicate] has been checked and adjusted if necessary.
This prompt ensures that the generated questions are grammatically correct, contextually appropriate, and free from potential underscoring issues, thereby reducing the risk of misinterpretation in subsequent AI processing stages. Using this methodology, we created 10,000 questions with balanced categories and reasoning types for testing purposes.
A.2 Response Generation and Confidence Measurement
We employed APIs on representative closed-source LLMs (GPT-3.5, GPT-4o, and GPT-o1) and open-source LLMs (Llama 3.1 8b and Llama 3.2 3b) to answer these questions independently, with temperature set to 0. We developed a prompt that not only elicits answers but also directs the AI to evaluate its own confidence levels regarding answer correctness, reasoning process, and facts used. The prompt used for generating these confidence-aware responses is as follows:
Please answer the following yes/no question ’[question]’. This question has only one correct answer. Follow these steps:
  1. 1.
    Think through the question step-by-step, employing a human-like reasoning process.
  2. 2.
    Pick the answer that you think is correct and begin with:
    • “The answer to the question is yes. … (reason)”
    • “The answer to the question is no. … (reason)”
    Even if you are unsure about the answer, pick the one that you think is more likely correct, and give your reasons.
  3. 3.
    Explain your reasoning process in detail.
  4. 4.
    List the key pieces of knowledge used in your reasoning, presented as declarative sentences and enumerated.
  5. 5.
    After providing your answer, evaluate your response in three aspects:
    • What is your estimate of the probability (in percent) that your answer is correct?
    • What is your estimate of the probability (in percent) that the facts underlying your answer are correct?
    • What is your estimate of the probability (in percent) that the reasoning underlying your answer is correct?
A.3 Similarity Measurement
A.3.1 Methodology Framework
The similarity score calculation methodology compares two sets of triples for each question ID: evidence triples and response triples (extracted from the model’s answer). The calculation follows a three-step process.
A.3.2 Facts Similarity Algorithm
Step 1: Subject Matching.
For each evidence triple, we identify relevant response triples based on subject matching criteria:
  • For subjects containing 1-2 words: Match is established if the response subject contains at least one identical word
  • For subjects containing 3+ words: Match is established if the response subject contains at least two identical words
Step 2: Object Similarity Calculation.
After identifying relevant subjects, we calculate object similarities using the following process:
Case 1:
For evidence triples with shared subjects:
  • Given evidence triples $[(A,\_,B),(A,\_,C)]$ and response triples $[(A,\_,B^{\prime}),(A,\_,C^{\prime}),(A,\_,D)]$
  • Calculate similarities: $sim(B,B^{\prime})$, $sim(C,C^{\prime})$
  • For non-direct matches (e.g., object $D$), calculate $sim(D,B)$ and $sim(D,C)$. Keep the higher similarity score for the final calculation
  • Store all similarity scores in $similarity\_list$
Case 2:
For evidence triples with different subjects:
  • Given evidence triples $[(A,\_,B),(C,\_,D)]$ and response triples $[(A,\_,B^{\prime}),(C,\_,E)]$
  • Calculate and store similarities: $sim(B,B^{\prime})$, $sim(D,E)$
Step 3: Final Score Calculation
We propose two scoring mechanisms to evaluate LLM’s response quality from different perspectives:
(1) Average Similarity Approach
This approach considers all generated content from the LLM, including potentially irrelevant information:
  • Store all calculated similarity scores
  • Final fact score = average of all similarity scores
This method provides a comprehensive evaluation of the LLM’s output, including its tendency to generate extraneous information.
(2) Maximum Similarity Approach
This approach focuses on the LLM’s best attempts to match the evidence:
  • For each evidence triple, retain only the highest similarity score among all matching response triples
  • Final fact score = average of these maximum similarity scores
This method evaluates the LLM’s ability to generate the most relevant and accurate information, regardless of any additional content provided.
A.3.3 Reasoning Similarity Algorithm
Step 1: Subject Matching.
For each evidence triple, we first assess subject matching with response triples. Response triples with non-matching subjects are excluded from calculation.
Step 2: Predicate and Object Similarity.
After subject matching, we evaluate predicate and object similarities according to the following cases:
Case 1: Similar Predicates
  • When predicates are similar, use the predicate similarity score as the triple’s similarity score
  • This applies regardless of object similarity
Case 2: Dissimilar Predicates
  • When predicates are not similar:
    • If objects are similar: Use predicate similarity as the triple’s score
    • If objects are not similar: Exclude triple from analysis
Step 3: Final Score Calculation.
  • Calculate similarity scores for each evidence triple using the above criteria
  • The final reasoning similarity score is the average of all valid triple similarity scores
Appendix B The Human Benchmark
B.1 Question Selection
To establish a human performance benchmark for comparison with LLM results, we selected a subset of 2,000 questions from our main dataset. The selection process involved careful manual verification to ensure that each question was unambiguous and that incorrect answers would stem from knowledge gaps or reasoning errors rather than linguistic confusion.
B.2 Experimental Protocol
The data was collected through Prolific’s online platform in January, 2025. Prior to the main task, participants received detailed instructions about the incentive mechanism, followed by a three-question comprehension test to verify their understanding of the procedure. Only participants who successfully passed this screening proceeded to the main experimental phase.
The core experimental task comprised ten independent rounds. In each round, participants responded to a binary-choice questions (requiring "Yes" or "No" responses) randomly selected from our validated question pool. Following each response, participants reported their confidence using a continuous scale ranging from 0 to 100, representing their subjective probability assessment of answer correctness.
B.3 Incentive Scheme
The compensation structure was as follows: Participants who failed the comprehension check received only the \$0.50 show-up fee and were excluded from the main task; If they complete the experiment, they would get \$ 2 as the base payment; Beyond this, they would get up to \$ 3 as the performance bonus based on the incentive mechanism following danz2022belief.
After each question, participants reported their confidence level, which determined their potential reward through the following mechanism:
For a randomly selected question, let $X\sim U[0,1]$ be a random draw and $c$ be the participant’s reported confidence: if $X<c$: Participant receives \$5 if their answer is correct; otherwise participant receives \$5 with probability $X$. This mechanism ensures that reporting one’s true confidence level is the optimal strategy so as to elicit truthful confidence reports.
B.4 Summary Statistics
For our primary analysis, we define two samples: one is the full sample (N=588), where all participants who completed the experiment; the other one is the baseline sample, where participants who complete the experiment and reported confidence $\geq 50\%$ for more than five out of ten questions (N=545).
Table LABEL:tab:002_benchmark_comparison_summary presents human performance by two measures across 10 rounds. "Diff." represents the difference between baseline sample and full sample and standard deviations are reported in parentheses. It shows no significant differences in performance patterns between the full and baseline samples, suggesting that our baseline criterion effectively identifies participants who engaged consistently with the confidence reporting task without introducing selection bias. In the main text, we use baseline samples to work as the benchmark of LLMs performance.
Table S1: Human Benchmark Performance Across Rounds
Accuracy Confidence
Round Baseline All Diff. Baseline All Diff.
1 0.63 0.63 -0.000 0.68 0.66 0.012
(0.212) (0.220)
2 0.66 0.66 0.006 0.67 0.65 0.018
(0.225) (0.235)
3 0.66 0.66 -0.002 0.69 0.68 0.014
(0.214) (0.224)
4 0.68 0.67 0.007 0.69 0.67 0.015
(0.205) (0.214)
5 0.65 0.65 -0.003 0.69 0.67 0.015
(0.214) (0.222)
6 0.66 0.66 0.008 0.68 0.67 0.016
(0.220) (0.229) 1
7 0.67 0.67 0.003 0.69 0.68 0.015
(0.212) (0.222)
8 0.64 0.64 0.003 0.67 0.66 0.014
(0.210) (0.218)
9 0.64 0.64 0.006 0.68 0.66 0.016
(0.214) (0.223)
10 0.66 0.66 0.002 0.69 0.67 0.015
(0.220) (0.229)
Accordingly, Table S2 presents a comparative analysis of confidence metrics between LLMs and human participants when responding to the same set of questions.
Table S2: Summary of Accuracy, Confidence, and Bias
Fraction of
correct answers
Confidence
in answer
Bias $N$
GPT 3.5 0.35 (0.476) 0.95 (0.043) 0.60 (0.473) 2000
GPT 4o 0.68 (0.465) 0.94 (0.056) 0.26 (0.451) 2000
GPT o1 0.81 (0.395) 0.96 (0.048) 0.15 (0.384) 2000
Llama 3.1 0.62 (0.486) 0.85 (0.113) 0.24 (0.461) 2000
Llama 3.2 0.62 (0.485) 0.95 (0.075) 0.33 (0.486) 2000
Human Baseline 0.66 (0.474) 0.70 (0.203) 0.04 (0.473) 5220
  • Notes: SDs are in parentheses. All biases are significantly different from zero ($p<0.0001$)
Appendix C The LLM-Exposure Experiment
C.1 Experimental Procotol
To examine how exposure to LLM outputs affects human decision-making and confidence calibration, we conducted a randomized controlled experiment with variations of exposure to LLM responses and their associated confidence levels.
We recruited participants through Prolific for a between-subjects experiment with three treatment conditions. The experiment proceeded in two stages. In Stage 1, all participants answered 10 questions randomly selected from the same pool as the human benchmark of 2,000 YES/NO reasoning questions and provided confidence assessments for each answer. In stage 2, participants were randomly assigned at the individual level to one of three experimental conditions:
  1. 1.
    Control: Participants were presented with their Stage 1 responses as defaults and given the opportunity to revise their answers and confidence levels without additional information.
  2. 2.
    LLM answer: Participants viewed the LLM’s answer (using either GPT-4o, GPT-o1, or Llama 3.1) before providing their revised answers and confidence assessments in the following format:
    [GPT-4o / GPT-o1 / Llama 3.1] answered [Yes/No] to this question.
  3. 3.
    LLM answer with confidence: Participants viewed both the LLM’s answer and its stated confidence level before providing their revised answers and confidence assessments:
    [GPT-4o / GPT-o1 / Llama 3.1] answered [Yes/No] to this question.
    [GPT-4o / GPT-o1 / Llama 3.1] indicated that its answer is correct with probability [x%].
In each experimental round, the specific LLM model presented to participants was randomly selected from among GPT-4o, GPT-o1, and Llama 3.1. These models were selected based on their higher accuracy compared to other available models.
We used the same incentive structure shown in section B.3. To ensure comprehension of the experimental procedures, participants who completed an example question with confidence assessment and answered three comprehension check questions about the instructions and incentive mechanism can enter the main part of the experiment.
C.2 Summary Statistics
There are 1364 participants who completed the experiment. We established our baseline sample by excluding participants who reported confidence exceeding 50% for more than five of the ten questions (N=1161). Table LABEL:tab:002_comparison_summary presents accuracy rates and confidence levels from Stage 1, showing mean values with standard deviations in parentheses across all rounds. As shown in the table, no significant differences in accuracy were observed between the baseline and full sample conditions. All subsequent analyses are based on the baseline data.
Table S3: Human Performance Across Rounds
Accuracy Confidence
Round Baseline All Diff. Baseline All Diff.
1 0.62 0.61 0.008 0.66 0.62 0.035
(0.212) (0.232)
2 0.64 0.62 0.016 0.66 0.63 0.038
(0.215) (0.235)
3 0.62 0.62 -0.002 0.68 0.64 0.035
(0.206) (0.228)
4 0.63 0.62 0.015 0.67 0.64 0.039
(0.200) (0.228)
5 0.65 0.64 0.008 0.68 0.64 0.041
(0.204) (0.229)
6 0.64 0.62 0.017 0.68 0.64 0.043
(0.201) (0.226)
7 0.63 0.63 0.003 0.68 0.64 0.039
(0.193) (0.222)
8 0.63 0.63 0.007 0.69 0.65 0.037
(0.196) (0.220)
9 0.64 0.64 0.009 0.69 0.65 0.038
(0.203) (0.225)
10 0.64 0.63 0.008 0.69 0.65 0.041
(0.197) (0.223)
Table LABEL:tab:002_human_exp_summary_stat presents summary statistics of pre-intervention accuracy and confidence across experimental conditions. We conducted one-way ANOVA tests on the first round baseline data to examine potential pre-existing differences among the experimental groups. For pre-intervention correctness, we found no significant differences among the three groups ($F(2,1158)=0.036$, $p=0.964$). Similarly, for pre-intervention confidence, no significant differences were observed ($F(2,1158)=0.609$, $p=0.544$). These results confirm successful randomization, with all groups demonstrating comparable baseline performance before experimental interventions were introduced.
Table S4: Accuracy Rates and Confidence Levels by Group Assignment
Accuracy Rate Confidence Level
Pre-test Post-test Pre-test Post-test $N$
Control 0.648 0.652 0.681 0.692 3960
(0.203) (0.200)
LLM answer 0.630 0.690 0.667 0.777 3860
(0.200) (0.185)
LLM answer+ confidence 0.624 0.699 0.685 0.845 3790
(0.207) (0.159)
Having established baseline equivalence across experimental conditions, we next examine participants’ behavioral patterns during task completion. Figure S1 illustrates the relationship between participants’ response times and accuracy rates across three experimental conditions. Response times were winsorized at the 95th percentile (70 seconds) and aggregated into 5-second intervals. The size of data points reflects the relative frequency of observations within each bin. All three conditions demonstrate comparable patterns, with accuracy rates generally fluctuating between 60% and 75% across the response time spectrum. When analyzing all 10 rounds collectively, the LLM answer with confidence group showed significantly faster pre-intervention response times compared to the other conditions (F(2, 11560) = 35.61, p < 0.001). However, our analysis of just the first round data revealed no significant differences in response times between groups (F(2, 1154) = 0.855, p = 0.425), again confirming successful randomization at baseline. This suggests that the observed response time differences emerged gradually over subsequent rounds rather than being present initially, indicating that repeated exposure to LLM answers and confidence information may progressively accelerate the answering process.
Figure S1: Average Accuracy Rate and Response Time across Experimental Conditions
While response time patterns reveal behavioral changes in how participants approached the tasks, our analysis of confidence calibration provides further insight into how LLM exposure influenced participants’ cognitive assessments of their own performance. Figure S2 reveals significant changes in confidence distributions between baseline (outlined black bars) and intervention stages (blue filled bars) across the three experimental conditions. In both treatment conditions (LLM answer and LLM answer with confidence), we observe a pronounced rightward shift from lower to higher confidence ranges following intervention. This shift is particularly notable in the group with exposure to both LLM-generated answers and confidence, in which it appears to substantially increase participants’ self-reported confidence levels.
Figure S2: Confidence Distribution (Baseline and Intervention Stage)
C.3 Subsample Analysis of Treatment Effects
We conduct a subsample analysis to investigate heterogeneous treatment effects based on confidence levels. We categorize observations along two dimensions: LLM confidence and human baseline confidence. For LLM confidence, we define "high" as instances where the selected LLM’s confidence exceeds 0.9, and "low" otherwise. For participants in the control group (who received no LLM assistance), we use the average confidence level across the three LLMs as the benchmark. For human baseline confidence, we classify observations as "high" when a participant’s confidence at the baseline stage exceeds the sample median, and "low" otherwise.
This classification yields four distinct subgroups: (1) low LLM confidence, low human confidence; (2) low LLM confidence, high human confidence; (3) high LLM confidence, low human confidence; and (4) high LLM confidence, high human confidence.
Table S5 presents summary statistics for these four subgroups. Panel A shows the distribution of observations, which is not uniform across categories, with the largest proportion (33.9%) in the high LLM confidence, high human confidence category. Panels B and C report human accuracy and bias at baseline, respectively, while Panels D and E present LLM accuracy and bias. Notably, LLM accuracy (Panel D) is substantially higher when LLM confidence is high (76.5-79.8%) compared to when it is low (50.1-51.8%), indicating that LLM confidence is a reliable predictor of accuracy.
Table S5: Combined Summary of LLM and Human Performance
Human Confidence
LLM Confidence Low High
Panel A: Observation Percentages
Low 19.3% 16.2%
High 29.4% 33.9%
Panel B: Human Accuracy
Low 0.543 (0.498) 0.677 (0.468)
High 0.545 (0.498) 0.745 (0.436)
Panel C: Human Bias
Low 0.493 (0.097) 0.37 (0.325)
High 0.49 (0.098) 0.314 (0.325)
Panel D: LLM Accuracy
Low 0.501 (0.457) 0.518 (0.446)
High 0.765 (0.401) 0.798 (0.376)
Panel E: LLM Bias
Low 0.467 (0.324) 0.445 (0.317)
High 0.247 (0.368) 0.215 (0.346)
Table S6 presents the correlation matrix between LLM accuracy, LLM confidence, baseline human accuracy, and baseline human confidence. A key observation is that the correlation between LLM confidence and human baseline confidence is quite modest (0.082), suggesting that LLM and humans tend to be uncertain about different questions. This low correlation has important welfare implications: individuals who are uncertain about an answer typically receive LLM inputs of average quality, which generally exceed the accuracy of humans who are uncertain at baseline. If uncertainty were more strongly correlated between humans and LLMs, we would expect LLM assistance to yield smaller accuracy improvements for low-confidence humans while potentially increasing bias.
Table S6: Correlation Matrix of LLM and Human Baseline Performance
LLM Accuracy LLM Confidence Baseline Accuracy Baseline Confidence
LLM Accuracy
LLM Confidence 0.289∗∗∗
Baseline Accuracy 0.182∗∗∗ 0.034∗∗∗
Baseline Confidence 0.061∗∗∗ 0.082∗∗∗ 0.214∗∗∗
  • Notes: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05.$
Table S7 presents the treatment effects across the four subgroups. The reference category is the group with high LLM confidence and high human baseline confidence. With the main treatment effects maintaining the positive impacts on accuracy, we observe significant negative interaction effects on accuracy (-8.1 to -8.9 percentage points) and substantial positive effects on bias (+20.8 to +22.4 percentage points) when exposed to LLM input for the low LLM confidence, low human confidence subgroup. This suggests that when both the human and LLM are uncertain, providing LLM input can be counterproductive.
Table S7: Subsample Treatment Effects across Experimental Conditions
$\Delta$Accuracy $\Delta$Bias
LLM Answer 0.051∗∗∗ -0.020
(0.011) (0.012)
LLM Answer+Confidence 0.056∗∗∗ 0.005
(0.012) (0.012)
Low LLM conf, Low human conf -0.010 0.025∗∗
(0.007) (0.008)
High LLM conf, Low human conf 0.002 0.015
(0.006) (0.007)
Low LLM conf, High human conf -0.002 0.001
(0.005) (0.005)
LLM Answer $\times$ Low LLM conf, Low human conf -0.089∗∗ 0.208∗∗∗
(0.027) (0.029)
LLM Answer+Confidence $\times$ Low LLM conf, Low human conf -0.081∗∗ 0.224∗∗∗
(0.031) (0.032)
LLM Answer $\times$ High LLM conf, Low human conf 0.100∗∗∗ 0.026
(0.021) (0.023)
LLM Answer+Confidence $\times$ High LLM conf, Low human conf 0.156∗∗∗ 0.080∗∗∗
(0.022) (0.022)
LLM Answer $\times$ Low LLM conf, High human conf -0.100∗∗∗ 0.091∗∗∗
(0.021) (0.023)
LLM Answer+Confidence $\times$ Low LLM conf, High human conf -0.095∗∗∗ 0.042
(0.025) (0.026)
Constant 0.005 -0.0002
(0.002) (0.003)
Observations 11,477 11,477
R2 0.03673 0.03143
  • Notes:${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$. Standard errors in parentheses are clustered at the player and question levels. The reference category is the group with high LLM confidence and high human baseline confidence.
C.4 Gender Gap in Overconfidence
Our analysis reveals gender differences in confidence that align with well-documented patterns in prior literature (see bordalo2019beliefshong and exley2022gender). Table LABEL:tab:008_gender_summary presents a comparative analysis of accuracy and confidence measures by gender across all experimental conditions, including baseline values, post-exposure values, and the magnitude of changes between stages.
The data demonstrate a consistent gender gap in confidence levels but not in accuracy. Across all experimental groups, male participants report significantly higher baseline confidence than female participants. These confidence differences exist despite comparable accuracy levels between genders in most conditions.
In response to LLM assistance, we observe notable gender differences in adaptability. In the LLM Answer group, women show larger confidence increases after exposure to LLM outputs compared to men (12.1 vs. 10.0 percentage points). In the LLM Answer with Confidence group, women demonstrate substantially larger accuracy improvements (9.1 percentage points) compared to men (6.0 percentage points), nearly eliminating the initial accuracy gap. The Control group shows minimal changes for both genders, confirming that observed effects in treatment groups are attributable to LLM assistance.
Table S8: Baseline Sample Summary Table by Gender and Groups
Measure Female Male Difference p.value
LLM Answer Group
     Baseline Accuracy 0.635 0.625 -0.009 0.544
     Post-exposure Accuracy 0.693 0.685 -0.008 0.598
     Accuracy Change 0.058 0.060 0.002 0.921
     Baseline Confidence 0.642 0.688 0.046 0.000
     Post-exposure Confidence 0.763 0.788 0.026 0.000
     Confidence Change 0.121 0.100 -0.020 0.000
LLM Answer with Confidence Group
     Baseline Accuracy 0.606 0.641 0.035 0.025
     Post-exposure Accuracy 0.696 0.701 0.005 0.742
     Accuracy Change 0.091 0.060 -0.031 0.074
     Baseline Confidence 0.669 0.700 0.031 0.000
     Post-exposure Confidence 0.829 0.859 0.030 0.000
     Confidence Change 0.160 0.159 -0.001 0.892
Control Group
     Baseline Accuracy 0.654 0.644 -0.010 0.491
     Post-exposure Accuracy 0.658 0.646 -0.012 0.411
     Accuracy Change 0.004 0.002 -0.002 0.682
     Baseline Confidence 0.670 0.689 0.018 0.005
     Post-exposure Confidence 0.683 0.700 0.017 0.007
     Confidence Change 0.013 0.011 -0.001 0.589
These results suggest that LLM assistance, particularly when accompanied by explicit confidence information, may help reduce gender-based differences in decision-making performance. The greater responsiveness of female participants to LLM inputs represents a promising avenue for using AI systems to potentially mitigate pre-existing confidence gaps.
Appendix D The Results in the Main Text
D.1 Table 1: Accuracy, confidence, and bias across models
D.1.1 Description of Table 1
Table 1 presents descriptive statistics for key variables across five Large Language Models (LLMs), based on their performance on 10,000 reasoning problems.There are a small proportion of missing values resulting from either premature termination of LLM outputs or instances where models expressed confidence in textual rather than numerical format. For each measure, we report both the mean value and its standard deviation (in parentheses).
Section (a) reports the fraction of correct answers, representing each model’s accuracy rate across the full sample. The results show substantial variation in performance across models, indicating meaningful differences in reasoning capabilities across models.
Section (b) presents self-reported confidence in the overall answer correctness ($\tilde{p}$). Notably, all models exhibit consistently high confidence, with most values exceeding 0.90, despite considerable variations in their actual performance.
Section (c) quantifies bias using difference between self-reported confidence ($\tilde{p}$) and actual accuracy. There is substantial positive bias across all models, indicating systematic overconfidence. Table S9 test whether bias are significantly different from zero. Panel A shows the estimates using self-reported confidence, while Panel B presents the results using our derived confidence measure ($\widetilde{p}_{F\cdot R}$) that combines the model’s confidence in facts and reasoning through multiplication ($p_{F}\cdot p_{R}$). We constructed this measure to address potential conjunction fallacies in confidence assessment, providing a more conservative estimate of overall confidence. The derived measure consistently yields lower values than the direct self-reported confidence in answers, suggesting that decomposing confidence assessment into constituent components may help mitigate overconfidence. The coefficient estimates (labeled "Overconfidence") represent the average bias for each model, with robust standard errors reported in parentheses.
The results provide strong statistical evidence of systematic overconfidence across all models regardless of whether we measure it using self-reported or derived confidence. All coefficients in these two subtables are positive and highly statistically significant ($p<0.001$), indicating that every model systematically overestimates its probability of being correct.
Table S9: Overconfidence Tests in Self-Reported and Derived Confidence
Panel A: Self-Reported Confidence
gpt 3.5 gpt 4o gpt o1 Llama 3.1 Llama 3.2
Overconfidence 0.59 0.30 0.22 0.23 0.33
(0.005) (0.005) (0.004) (0.005) (0.005)
$N$ 9,795 9,973 9,990 9,750 9,910
Panel B: Derived Confidence
gpt 3.5 gpt 4o gpt o1 Llama 3.1 Llama 3.2
Overconfidence 0.52 0.24 0.17 0.16 0.24
(0.005) (0.005) (0.004) (0.005) (0.005)
$N$ 9,779 9,895 9,974 9,614 9,854
D.2 Figure 1: LLM Confidence and Accuracy
D.2.1 Description of Figure 1
Figure 1 presents the visualization of the relationship between confidence and accuracy across LLM models and the human benchmark.
In all panels of Figure 1, the horizontal axis represents the self-reported confidence levels and the vertical axis represents the mean accuracy rate conditional at each confidence level. The 45-degree line represents perfect calibration, where confidence exactly matches accuracy. Points below this line indicate overconfidence, while points above would indicate underconfidence. The size of each plotted point is proportional to the number of observations at that confidence level.
Panel A focuses on the GPT family of models (GPT 3.5, GPT 4o, and GPT o1), while Panel B presents the corresponding analysis for Llama models (Llama 3.1 8B and Llama 3.2 3B), both with 10,000 observations. Further, Panel C provides a comparative analysis between advanced LLM models (GPT 4o, GPT o1, and Llama 3.1 8B) and human performance, based on the human benchmark subset of 2,000 questions.
The fitted lines represent weighted regression coefficients, where observations are weighted by their frequency. The shaded areas around each fitted line represent 95% confidence bands, indicating the precision of our estimates. All slope coefficients (coefficients correspond to Panel A and B are detailed in Table 2 and those for Panel C are detailed in Table S11) are statistically significant ($p<0.001$) across models and comparison groups, suggesting robust relationships between confidence and accuracy, albeit with systematic overconfidence.
D.2.2 Comparison of Confidence Measures
To examine whether the patterns observed with self-reported confidence persist when using a measure that potentially mitigates conjunction fallacy, we also conducted parallel analyses on LLMs using our derived confidence measure ($\widetilde{p}_{F\cdot R}$) in Table S10.
Table S10: Derived Confidence gradients in LLM accuracy
gpt 3.5 gpt 4o gpt o1 llama3.1 llama3.2
Confidence 0.30 1.54 1.64 0.90 0.59
gradient (0.06) (0.06) (0.06) (0.04) (0.04)
Predicted accuracy 0.39 0.82 0.89 0.82 0.70
if $\widetilde{p}_{F\cdot R}=1$ (0.01) (0.01) (0.01) (0.01) (0.01)
$N$ 9,779 9,895 9,974 9,614 9,854
R2 0.003 0.06 0.07 0.05 0.02
  • Notes: Dependent variable is correct answer (= 1). Confidence levels are normalized by subtracting 1 from the original confidence scores, so the constant term represents accuracy at 100% confidence ($\tilde{p}=1$). Regression is weighted by the frequency of confidence levels. All coefficients are significant at p<0.01.
The analysis reveals similar qualitative patterns across both GPT and Llama model families, as illustrated in Panels A and B of Figure S3. However, Panel C reveals an important nuance in the relationship between self-reported and derived confidence measures. While the two measures are highly correlated, self-reported confidence consistently exceeds the derived measure, providing evidence of conjunction bias in model self-assessment. Notably, the difference between self-reported and derived confidence narrows at higher confidence levels, suggesting that models’ assessments become more internally consistent—and potentially more rational—when they are most confident in their answers.
Figure S3: LLM Derived Confidence and Accuracy
D.3 Table 2: The confidence gradient in accuracy
D.3.1 Description of Table 2
Table 2 presents the regression analysis quantifying the relationship between accuracy and confidence across different models using the full sample of questions. For each source, we estimate:
$\mathbf{1}_{\{correct,i\}}=\beta_{0}+\beta_{1}(Confidence_{i}-1)+\epsilon_{i}$
where $\mathbf{1}_{\{correct,i\}}$ is an indicator equal to 1 if answer $i$ is correct, and $Confidence_{i}$ represents either self-reported confidence ($\widetilde{p}$) for all sources or derived confidence ($\widetilde{p}_{F\cdot R}$) for models. The confidence measures are normalized by subtracting 1, so that $\beta_{0}$ directly represents predicted accuracy at full confidence.
The confidence gradient ($\beta_{1}$) quantifies how accuracy changes with confidence levels, corresponding to the values of the fitting lines of Panel A and B in Figure 1. The predicted accuracy at full confidence ($\beta_{0}$) reveals systematic overconfidence. The regression estimates are weighted by the frequency of observations at each confidence level and heteroskedasticity-consistent standard errors (HC0) are reported in parentheses. The results reveal substantial heterogeneity across models. More advanced models (such as GPT 4o and GPT o1) exhibit steeper confidence gradients, indicating a stronger relationship between their confidence assessments and actual performance. However, even these models show significant overconfidence at full confidence levels, as evidenced by their intercept terms being consistently below 1.
D.3.2 Confidence Gradients on Human Benchmark Subset
We extend our analysis by estimating confidence gradients for the five LLM models and human benchmark using the same set of 2000 questions. These estimates correspond to the fitted lines displayed in Panels C of Figures 1.
Confidence levels are normalized by subtracting 1 from the original confidence scores, so the constant term represents accuracy at 100% confidence ($\tilde{p}=1$). These regression estimates are weighted by the frequency of observations at each confidence level. For human participants, standard errors are clustered at both individual and question levels. All reported coefficients are statistically significant at p < 0.01.
Table S11: The confidence gradient in accuracy (Human Baseline Subsample)
human gpt_3.5 gpt_4o gpt_o1 llama3.1_8b llama3.2_3b
Confidence 0.52 1.49 2.65 2.61 1.55 0.42
gradient (0.04) (0.26) (0.19) (0.28) (0.09) (0.15)
Predicted accuracy 0.81 0.42 0.84 0.92 0.84 0.64
if $\tilde{p}=1$ (0.01) (0.02) (0.01) (0.01) (0.02) (0.01)
$N$ 5,220 1,933 1,994 1,997 1,958 1,984
$R^{2}$ 0.05 0.02 0.09 0.08 0.11 0.003
D.4 Figure 2: Treatment Effects on Changes in Accuracy and Bias
D.4.1 Description of Figure 2
The key variables of interest in the LLM-exposure experiment are the differences in accuracy and bias (measured by the gap between self-reported confidence and actual correctness) between Pre-treatment and post-treatment measurements.Table S12 presents regression results across treatment conditions, where $\Delta$accuracy and $\Delta$bias represent the differences between baseline and intervention stages for each respective measure.
Table S12: Treatment Effects across Experimental Conditions
$\Delta$Accuracy $\Delta$Bias
LLM Answer 0.056∗∗∗ 0.042∗∗∗
(0.010) (0.011)
LLM Answer + Confidence 0.071∗∗∗ 0.076∗∗∗
(0.010) (0.011)
Constant 0.003 0.009∗∗∗
(0.002) (0.003)
Observations 11,610 11,610
R2 0.00530 0.00522
  • Notes: ${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$. Standard errors in parentheses are clustered at the player and question levels.
The regression results reveal substantial treatment effects across outcome variables. Both experimental conditions significantly improved accuracy rates, with LLM Answer+Confidence Group showing a more pronounced effect (7.1 percentage points) compared to the LLM Answer Group (5.6 percentage points) but there is no statistically significant difference in accuracy improvement between these two treatment groups ($\chi^{2}(1)=1.55$, $p=0.213$). The full treatment effects for each experimental condition are illustrated in the Panel A of Figure 2 in the main text.
Notably, both treatments significantly increased participants’ bias with the LLM Answer+Confidence Group exhibiting a stronger effect (7.6 percentage points) compared to the LLM Answer Group (4.2 percentage points) and this difference in bias induction between treatment groups was statistically significant ($\chi^{2}(1)=6.46$, $p=0.011$). The full treatment effects on bias are displayed in the Panel B of Figure 2 in the main text.
D.5 Table 3: Subsample Analysis of LLM Exposure Experiment
In table 3, we examine heterogeneity across two dimensions: (1) how the presence of LLM confidence information influences participants’ judgment and decision-making, (2) how participants’ initial confidence levels moderate the effects of LLM assistance. This analysis provides insights into the contexts and populations for whom LLM assistance offers the greatest benefit.
D.5.1 Conditional on LLM Confidence
To investigate how LLM confidence levels moderate treatment effects, we constructed LLM confidence measures that varied by experimental condition. For the control group, we used the average confidence level from GPT-4o, GPT-o1, and Llama 3.1 on each question. For treatment groups (both LLM answer and LLM answer with confidence), we used the confidence level expressed by the specific model that participants encountered for each question.
To examine whether high LLM confidence levels distinctly influence treatment effects, columns 1 and 2 of Table 3 present regression results using the subsample with LLM confidence above 0.8 (on a 0-1 scale). We define high confidence as a binary variable indicating LLM confidence scores above 0.9. The combined treatment effects with 95% confidence intervals for each experimental condition are illustrated in the first part of Panel A and Panel B of Figure S4 separately.
Panel A: Heterogeneous Treatment Effects on Change in Accuracy
Panel B: Heterogeneous Treatment Effects on Change in Bias
Figure S4: Illustration of Subsample Analysis
The findings reveal striking interaction effects. While the main treatment effects show minimal impact on accuracy changes, their interactions with high confidence demonstrate substantial effects. Specifically, when LLMs express high confidence (confidence ranging from 0.9 to 1), the LLM Answer treatment increases accuracy by 13.6 percentage points, while the LLM Answer with Confidence treatment produces an even larger improvement of 15.3 percentage points. Similarly, both treatments significantly reduce bias when coupled with high LLM confidence. The interaction between high confidence and the LLM Answer treatment reduces bias by 13.8 percentage points, while the interaction with the LLM Answer with Confidence treatment reduces bias by 11.9 percentage points. Such results show that the benefits of LLM assistance are concentrated in scenarios where the models express high confidence in their answers, suggesting that confidence signals contain valuable information that helps participants discriminate between reliable and unreliable AI outputs.
D.5.2 Conditional on Human Baseline Confidence
Understanding how initial user confidence mediates the effects of LLM assistance is also crucial for optimizing human-AI collaboration. Columns 3 and 4 of Table 3 present our heterogeneity analysis examining how participants’ baseline confidence levels (i.e., their initial confidence at pre-treatment stage) moderate the effects of LLM answers on accuracy and bias, where high confidence is defined as above median baseline confidence. Accordingly, combined treatment effects are illustrated in the second part of Panel A and Panel B in Figure S4 separately.
The results reveal substantial heterogeneity in treatment effects. Specifically, high-confidence participants show reductions of 6.4 percentage points in the accuracy benefit from LLM Answer and 9.1 percentage points in the benefit from LLM Answer+Confidence. The pattern is similar for bias reduction, with high-confidence participants showing reductions of 6.3 and 12.4 percentage points respectively in the treatment benefits.
These results suggest that LLM assistance provides the greatest value to users who lack confidence in their initial judgments, potentially because they are more receptive to incorporating AI suggestions. Conversely, highly confident users appear more resistant to revising their judgments based on LLM inputs, resulting in significantly smaller improvements from AI assistance.
D.5.3 Subsample Analysis: Continuous Version
To provide a more comprehensive picture of heterogeneous treatment effects and to serve as a robustness check for Table 3, Table S13 presents regression results examining how continuous measures of LLM confidence and human baseline confidence relate to changes in participant accuracy and bias across experimental conditions.
Table S13: Continuous Analysis of Treatment Effects Moderated by LLM and Human Confidence
Continuous LLM Confidence Baseline Human Confidence
$\Delta$ Accuracy $\Delta$ Bias $\Delta$ Accuracy $\Delta$ Bias
LLM Answer -0.595∗∗∗ 0.688∗∗∗ 0.086∗∗∗ 0.070∗∗∗
(0.101) (0.108) (0.014) (0.017)
LLM Answer+Confidence -0.555∗∗∗ 0.310∗∗ 0.114∗∗∗ 0.139∗∗∗
(0.098) (0.098) (0.016) (0.017)
LLM Confidence 0.053 -0.053
(0.036) (0.041)
LLM Answer 0.707∗∗∗ -0.702∗∗∗
$\times$ LLM Confidence (0.109) (0.115)
LLM Answer+Confidence 0.685∗∗∗ -0.256
$\times$ LLM Confidence (0.106) (0.106)
Baseline Confidence -0.010 -0.050∗∗∗
(0.012) (0.015)
LLM Answer -0.177∗∗∗ -0.174∗∗∗
$\times$ Baseline Confidence (0.040) (0.045)
LLM Answer+Confidence -0.233∗∗∗ -0.338∗∗∗
$\times$ Baseline Confidence (0.044) (0.047)
Constant -0.046 0.057 0.005 0.018∗∗∗
(0.033) (0.038) (0.004) (0.005)
Observations 11,477 11,477 11,610 11,610
R2 0.02181 0.01409 0.01256 0.02005
  • Notes: Models 1-2 include LLM confidence (mean-centered at 0.8), and Models 3-4 use respondents’ prior confidence as a moderator.
  • ${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$.
  • Standard errors in parentheses are clustered at the participant and question levels.
Here’s a more concise version:
For LLM confidence, our findings reveal significant patterns. While main treatment effects on accuracy are slightly negative, these effects are substantially moderated by LLM confidence levels. The interaction between LLM confidence and both treatment conditions shows strong positive effects on accuracy changes and corresponding negative effects on bias changes. This indicates that when LLMs express higher confidence, their answers produce greater improvements in participant accuracy, though bias reduction may be less pronounced when LLM confidence is explicitly shared.
When examining human baseline confidence, participants shown LLM answers experience accuracy improvements of 8.6 percentage points compared to controls, increasing to 11.9 percentage points when LLM confidence is also provided. Similar patterns emerge for bias reduction (7.0 and 13.9 percentage points respectively). However, the significant negative interaction coefficients demonstrate that participants with higher baseline confidence benefit substantially less from LLM assistance, with this effect being even more pronounced for bias reduction.
These continuous variable analyses confirm the patterns observed in our subsample analysis, providing robust evidence that LLM assistance is most beneficial when models express high confidence and when users initially express low confidence in their judgments.
D.6 Descriptive Statistics
D.6.1 Distribution of Question Types
Our question datasets comprises 10,000 questions balanced across domains and reasoning types. Table S14 shows the distribution of questions, which are composed of five different question types (composite, inverse, negation, temporal, and transitive reasoning) across ten knowledge domains. Each cell contains 200 questions, resulting in 1,000 questions per domain and 2,000 questions per question type.
Table S14: Distribution of Question Types Across Domains
Domain Composite Inverse Negation Temporal Transitive Total
Culture 200 200 200 200 200 1,000
Geography 200 200 200 200 200 1,000
Health 200 200 200 200 200 1,000
History 200 200 200 200 200 1,000
Math 200 200 200 200 200 1,000
Nature 200 200 200 200 200 1,000
People 200 200 200 200 200 1,000
Religion 200 200 200 200 200 1,000
Society 200 200 200 200 200 1,000
Technology 200 200 200 200 200 1,000
Total 2,000 2,000 2,000 2,000 2,000 10,000
D.6.2 Distribution of Confidence Measures
The figures below show the distribution of different confidence measures across all five LLM models in our analysis. In all figures, the x-axis represents confidence levels, and the y-axis shows the percentage of observations in each bin with the width of 0.05. Figure S5 displays the distribution of self-reported confidence in answer correctness; Figure S6 and Figure S7 show the distribution of confidence in factual knowledge and reasoning process, respectively; Figure S8 presents the distribution of our derived confidence measure (multiplication of confidence in facts and reasoning, with the range from 0 to 1).
The distributions reveal several notable patterns. First, all models exhibit a strong rightward skew in their confidence distributions, with a substantial mass of observations at very high confidence levels. This pattern is particularly pronounced for the GPT family of models. Second, the derived confidence measure (Figure S8) shows generally lower and more dispersed values compared to the direct self-reported confidence, consistent with our conjunction fallacy hypothesis. Third, there are systematic differences between model families, with the Llama models showing somewhat more dispersed confidence distributions compared to their GPT counterparts.
Figure S5: Distribution of Confidence in Answer Correctness across LLMs
Figure S6: Distribution of Confidence in Facts Correctness across LLMs
Figure S7: Distribution of Confidence in Reasoning Correctness across LLMs
Figure S8: Distribution of Derived Confidence in Correctness across LLMs
D.7 Similarity Measure Results
Table D.7 and Table D.7 present pairwise correlations between the main variables for GPT and llama family of models, respectively. Here, we consider two types of similarity scores: similarity scores in average and similarity score in maximum, both computed for reasoning and facts components separately. There are high correlations between average and maximum scores scross all models, suggesting that these alternative measures could capture the similar underlining patterns.
In both tables, the positive relationship between self-reported (derived) confidence and similarity measures show consistent patterns for both facts and reasoning components. The correlations coefficients between relf-reported (derived) confidence and similarity are highly significant across all the model, though the magnitudes vary between two families of models.
Table S15: Correlation Matrices: GPT Models
Panel A: GPT 3.5
Panel B: GPT 4o    Variables (1) (2) (3) (4) (5) (6) (7) (8) (9) 1. Correct Answer 1 2. Confidence in Answer 0.27*** 1 3. Derived Confidence 0.26*** 0.88*** 1 4. Confidence in Facts 0.21*** 0.76*** 0.91*** 1 5. Confidence in Reasoning 0.26*** 0.89*** 0.95*** 0.74*** 1 6. Average Fact Score 0.10*** 0.24*** 0.25*** 0.18*** 0.27*** 1 7. Average Reasoning Score 0.07*** 0.23*** 0.23*** 0.17*** 0.25*** 0.91*** 1 8. Maximum Fact Score 0.10*** 0.24*** 0.24*** 0.18*** 0.25*** 0.96*** 0.88*** 1 9. Maximum Reasoning Score 0.07*** 0.23*** 0.23*** 0.17*** 0.24*** 0.87*** 0.96*** 0.90*** 1    Panel C: GPT o1    Variables (1) (2) (3) (4) (5) (6) (7) (8) (9) 1. Correct Answer 1 2. Confidence in Answer 0.24*** 1 3. Derived Confidence 0.26*** 0.87*** 1 4. Confidence in Facts 0.23*** 0.81*** 0.97*** 1 5. Confidence in Reasoning 0.25*** 0.84*** 0.89*** 0.75*** 1 6. Average Fact Score 0.12*** 0.13*** 0.13*** 0.10*** 0.16*** 1 7. Average Reasoning Score 0.10*** 0.13*** 0.13*** 0.10*** 0.16*** 0.90*** 1 8. Maximum Fact Score 0.11*** 0.14*** 0.13*** 0.10*** 0.16*** 0.97*** 0.87*** 1 9. Maximum Reasoning Score 0.09*** 0.13*** 0.13*** 0.10*** 0.15*** 0.87*** 0.97*** 0.90*** 1
Table S16: Correlation Matrices: Llama Models
Panel B: Llama 3.2    Variables (1) (2) (3) (4) (5) (6) (7) (8) (9) 1. Correct Answer 1 2. Confidence in Answer 0.09*** 1 3. Derived Confidence 0.14*** 0.73*** 1 4. Confidence in Facts 0.09*** 0.53*** 0.78*** 1 5. Confidence in Reasoning 0.13*** 0.72*** 0.94*** 0.52*** 1 6. Average Fact Score 0.03** 0.16*** 0.14*** 0.09*** 0.15*** 1 7. Average Reasoning Score 0.03** 0.16*** 0.14*** 0.09*** 0.14*** 0.90*** 1 8. Maximum Fact Score 0.04*** 0.16*** 0.14*** 0.10*** 0.14*** 0.97*** 0.87*** 1 9. Maximum Reasoning Score 0.03*** 0.16*** 0.14*** 0.10*** 0.14*** 0.88*** 0.97*** 0.90*** 1
D.8 Replication Results
D.8.1 Descriptive Summary
To validate the reliability of our LLM-generated responses, we conducted four additional rounds of experiments with GPT-3.5 and GPT-4o in December 2024, followed by another round in February 2025, yielding a total of six experimental rounds for each model. Each round maintained identical prompt structures and problem sets (N = 10,000) while implementing randomized sequences of the question list to control for potential order effects. Table LABEL:tab:006_replication_summary provides a detailed summary of the correctness and confidence measures across all six rounds. The fraction of correct answers shows the accuracy rate, while the confidence measures show the model’s self-reported confidence in its answers, facts, and reasoning respectively. The mean confidence measure represents the average of all confidence metrics.
Table S17: Replication Results: Accuracy and Confidence Measures
a) Fraction of b) LLM confidence in … c) Derived
Round correct answers answer ($\tilde{p}$) facts ($p_{F}$) reasoning ($p_{R}$) confidence ($\tilde{p}_{F,R}$)
GPT 3.5
     1 0.35 0.94 0.96 0.91 0.87
(0.476) (0.045) (0.044) (0.061) (0.082)
     2 0.35 0.94 0.95 0.91 0.87
(0.476) (0.045) (0.044) (0.061) (0.081)
     3 0.35 0.94 0.95 0.91 0.87
(0.477) (0.045) (0.045) (0.061) (0.082)
     4 0.35 0.94 0.96 0.91 0.87
(0.477) (0.045) (0.043) (0.061) (0.082)
     5 0.35 0.94 0.96 0.91 0.87
(0.477) (0.045) (0.044) (0.061) (0.082)
     6 0.35 0.94 0.95 0.91 0.87
(0.476) (0.045) (0.045) (0.061) (0.082)
GPT 4o
     1 0.63 0.94 0.95 0.93 0.88
(0.482) (0.053) (0.044) (0.054) (0.083)
     2 0.65 0.93 0.94 0.92 0.87
(0.478) (0.058) (0.048) (0.055) (0.085)
     3 0.65 0.93 0.94 0.92 0.87
(0.478) (0.058) (0.049) (0.055) (0.085)
     4 0.64 0.93 0.94 0.92 0.87
(0.479) (0.059) (0.048) (0.054) (0.085)
     5 0.64 0.93 0.94 0.92 0.87
(0.479) (0.059) (0.049) (0.055) (0.086)
     6 0.65 0.93 0.94 0.93 0.87
(0.476) (0.055) (0.046) (0.052) (0.082)
For the the pairwise differences across rounds, Table LABEL:tab:007_gpt35_rounds_diffference and Table LABEL:tab:007_gpt4o_rounds_diffference present the difference between two rounds for each model with the average and standard deviation in parenthesis and also p-values. We also report the detailed distribution of response patterns for these two models in Figure S9. Each pattern is represented by a six-digit binary sequence, where ’1’ indicates a correct answer and ’0’ indicates an incorrect answer in a given round. For example, ’111111’ represents consistent correctness across all rounds, while ’000000’ represents consistent incorrectness. These pattern distributions reveals high consistency across rounds in binary responses, with GPT-3.5 and GPT-4o demonstrating consistency rates of 87.58% and 92.3%, respectively.
Panel A: GPT-3.5
Panel B: GPT-4o
Figure S9: Proportion of Correctness Combinations Across Six Rounds
Table S18: Differences in Accuracy and Confidence: GPT-3.5
Accuracy Confidence
Rounds Mean (SD) p-value Mean (SD) p-value
1_2 0.001 (0.238) 0.733 -0.000 (0.040) 0.810
1_3 0.002 (0.236) 0.439 -0.000 (0.039) 0.469
1_4 0.002 (0.239) 0.309 -0.000 (0.040) 0.291
1_5 0.002 (0.238) 0.395 -0.000 (0.039) 0.378
1_6 -0.002 (0.258) 0.346 0.000 (0.042) 0.998
2_3 0.001 (0.219) 0.611 -0.000 (0.039) 0.622
2_4 0.002 (0.216) 0.372 -0.000 (0.039) 0.397
2_5 0.001 (0.226) 0.822 -0.000 (0.040) 0.531
2_6 -0.003 (0.254) 0.263 0.000 (0.042) 0.809
3_4 0.001 (0.215) 0.572 -0.000 (0.039) 0.725
3_5 -0.001 (0.224) 0.751 -0.000 (0.040) 0.885
3_6 -0.005 (0.257) 0.076 0.000 (0.042) 0.490
4_5 -0.001 (0.217) 0.708 0.000 (0.039) 0.836
4_6 -0.005 (0.258) 0.065 0.000 (0.042) 0.311
5_6 -0.004 (0.257) 0.114 0.000 (0.042) 0.410
Table S19: Differences in Accuracy and Confidence: GPT-4o
Accuracy Confidence
Rounds Mean (SD) p-value Mean (SD) p-value
1_2 0.012 (0.213) $0$0.006 (0.037) 0.000
1_3 0.012 (0.216) $0$0.006 (0.037) 0.000
1_4 0.010 (0.215) $0$0.007 (0.036) 0.000
1_5 0.009 (0.211) $0$0.006 (0.039) 0.000
1_6 0.017 (0.238) $0$0.003 (0.038) 0.000
2_3 -0.000 (0.096) 0.917 -0.000 (0.029) 0.760
2_4 -0.002 (0.116) 0.143 -0.000 (0.031) 0.183
2_5 -0.002 (0.112) 0.033 -0.000 (0.030) 0.796
2_6 0.005 (0.189) 0.007 0.003 (0.035) 0.000
3_4 -0.002 (0.117) 0.170 -0.000 (0.028) 0.259
3_5 -0.002 (0.107) 0.032 0.000 (0.029) 0.975
3_6 0.005 (0.195) 0.008 0.003 (0.036) 0.000
4_5 -0.001 (0.114) 0.538 0.000 (0.030) 0.273
4_6 0.007 (0.199) 0.001 0.003 (0.037) 0.000
5_6 0.008 (0.196) 0.000 0.003 (0.037) 0.000
D.8.2 Regression Analysis of Response Persistence
To formally assess the stability of LLM responses across 5 rounds, we conduct three types of regression analyses for both GPT-3.5 and GPT-4o. Table S20 presents the results from these analyses.
Model 1 (columns 1-2) examines the persistence in response accuracy across rounds. The strong coefficients on previous accuracy (0.8830 for GPT-3.5 and 0.9458 for GPT-4o, both significant at the 0.1% level) indicate high consistency in whether responses are correct across rounds.
Model 2 (columns 3-4) analyzes how previous accuracy influences confidence levels in subsequent rounds. While statistically significant, the small coefficients (0.0113 for GPT-3.5 and 0.0305 for GPT-4o) suggest that previous accuracy has limited impact on future confidence levels. The high constant terms (0.9384 and 0.9384) indicate that both models maintain high baseline confidence levels regardless of previous performance.
Model 3 (columns 5-6) combines both effects by including both previous confidence and previous accuracy on current confidence. The results reveal strong persistence in confidence levels, particularly for GPT-4o (coefficient of 0.8365 compared to 0.6035 for GPT-3.5). Once controlling for previous confidence, the effect of previous accuracy becomes small, suggesting again that confidence patterns are largely independent of actual performance.
Across all specifications, round effects are minimal, with coefficients close to zero, which indicates that the sequence of questions does not systematically influence either accuracy or confidence.
Table S20: Response Persistence and Round Effects
Response Correctness Confidence Level Combined Analysis
GPT-3.5 GPT-4o GPT-3.5 GPT-4o GPT-3.5 GPT-4o
Previous Accuracy 0.8830*** 0.9458*** 0.0113*** 0.0305*** 0.0041*** 0.0046***
(0.0035) (0.0021) (0.0007) (0.0012) (0.0003) (0.0003)
Previous Confidence 0.6035*** 0.8365***
(0.0092) (0.0069)
Round 3 0.0004 -0.0112*** -0.0002 -0.0004* -0.0002 0.0051***
(0.0037) (0.0024) (0.0004) (0.0003) (0.0006) (0.0005)
Round 4 0.0006 -0.0127*** -0.0004* -0.0008* -0.0002 0.0048***
(0.0030) (0.0024) (0.0004) (0.0003) (0.0005) (0.0004)
Round 5 -0.0014 -0.0119*** -0.0003* -0.0004* -0.0000 0.0055***
(0.0031) (0.0023) (0.0004) (0.0003) (0.0005) (0.0004)
Round 6 -0.0049* -0.0036* -0.0000 0.0026*** 0.0002 0.0081***
(0.0034) (0.0030) (0.0004) (0.0004) (0.0005) (0.0005)
Question Order $0**$$0$$0$$0**$$0$0.0000*
(0.0000) (0.0000) (0.0000) (0.0000) (0.0000) (0.0000)
Constant 0.0360*** 0.0463*** 0.9384*** 0.9136*** 0.3720*** 0.1446***
(0.0029) (0.0029) (0.0006) (0.0012) (0.0087) (0.0064)
Num.Obs. 48588 49996 48986 49994 48983 49965
R2 0.781 0.898 0.015 0.065 0.372 0.711
  • Notes: ***Significant at the 1 percent level; **Significant at the 5 percent level; *Significant at the 10 percent level.
Appendix E The welfare effects of LLM exposure
This section describes the welfare analysis discussed in the main body of the paper.
E.1 The setup
Assume there is an investment project. The payoff from the project depend on (i) choosing the true state of the world (labeled $Y$ or $N$, in analogy to the questions we ask in our experiment), and (ii) an investment $e$ that increases the payoff from the project if the correct answer is chosen. Investment has a cost $c(e)$, with $c^{\prime}(e)>0$ and $c^{\prime\prime}(e)\geq 0$. If the correct answer is chosen, the payoff is $\pi e$. If the incorrect answer is chosen, the payoff is zero.
The has beliefs $\widetilde{p}=p+b$, where $p$ is the true probability that the correct state is chosen, and $b$ is an individual bias in assessing the accuracy of the judgment. We focus here on the case where $b>0$, i.e. the individual is overconfident in his response.
The individual thus chooses the state of the world that has $\widetilde{p}\geq 0.5$. His perceived utility from investing in the project is given by
$V(e;\widetilde{p})=(p+b)\cdot\pi\cdot e-c(e)$ (1)
The individual maximizes equation (1) by choosing effort such that
$\frac{\partial V}{\partial e}$ $=(p+b)\pi-c^{\prime}(e)=0$ (2)
i.e. he chooses effort such that the marginal cost of effort $c^{\prime}(e)$ equal the perceived marginal benefit $\pi(p+b)$. Optimal perceived effort, denoted $e(p+b)$ henceforth, is thus defined by
$e(p+b):c^{\prime}(e)=(p+b)\pi$ (3)
It is straightforward to show that $\frac{\partial e(p+b)}{\partial\widetilde{p}}>0$. We will use the shorthand notation $e^{\prime}()$ to refer to this derivate below. Notice that an overconfident individual overinests into the project, since he perceives the the marginal benefit to be $\pi(p+b)$ instead of the true $\pi p$.
We will at times use the special case of $c(e)=\exp(e\gamma)/\gamma$. In this case, $c^{\prime}(e)=\exp(e\gamma)$. Optimal effort is given by $e(p+b)=\frac{\ln(\pi(p+b))}{\gamma}$, and $e^{\prime}(p+b)=\frac{1}{\gamma(p+b)}$.
Following bernheim2019behavioral, the indivdiual’s true welfare is given by
$W(p,b)=p\pi e(p+b)-c(e(p+b))$ (4)
i.e. it evaluates the welfare using the objective probability $p$, but takes into account that the individual chooses effort acording to equation (3), based on his overconfident belief $\widetilde{p}$. This allows us to get a first glimpse at the welfare effects of changing $p$ and $b$.
In this model, exposure to LLMs potentially changes, both, $p$ and $b$. To derive the first two results we discuss in the text, we evaluate the first derivatives of $W$ with regard to these two parameters. The derivative of $W$ with respect to $b$ is given by
$\frac{\partial W}{\partial b}$ $=p\pi e^{\prime}(p+b)-c^{\prime}(e)\cdot e^{\prime}(p+b)$ (5)
$=-b\pi e^{\prime}<0$
where the simplications in the second line of the equation follow from observing that the indivdiual chooses $e$ such that $c^{\prime}(e)=\pi(p+b)$, and substituting this. Unsurprisingly, welfare decreases in $b$ unambiguously, reflecting the fact that a higher $b$ leads to higher over-investment.
The derivative of $W$ with respect to $p$ is given by
$\frac{\partial W}{\partial p}$ $=\pi e+p\pi e^{\prime}-c^{\prime}(e)$ (6)
$=\pi e-b\pi e^{\prime}$
Notice that, somewhat surprisingly, the derivative is not unambiguously positive if $b>0$ and $e^{\prime}>0$. That is, it is possible that welfare decreases if the objective proability of success increases. We discuss this effect in more detail below.
These derivatives allow us to state the first two results.
Result 1: if $e^{\prime}(p+b)=0$, then welfare unambiguously increases in $p$.
This is a very intuitive result. Since $e^{\prime}(p+b)=0$, effort does not change a function of $\widetilde{p}$. Thus, there is no cost from overinvestment, and payoffs unambiguously increase as $p$ increases.
Result 2: If $e^{\prime}(p+b)>0$, then welfare decreases in $p$ if
$\pi e<b\pi e^{\prime}$
In the case of an exponential cost function, this is the case if
$\ln(\pi(p+b))<\frac{b}{p+b}$ (7)
Intuitively speaking, the condition for this effect to occur is if $e$ is relatively low, but $e(p+b)$ is elastic. If $e(p+b)$ responds very strong to changes in $p$. In this case, an increase in $p$ has a direct on utility, raising payoffs by $\pi e$ per unit of $p$. However, a larger $p$ also increases investment. Recall however, that, the individual over-invests because of $b$. Thus, this increase creates increasingly expensive over-investment. If effort is elastic enough, this second effect can dominate the first and decrease overall welfare.
Equation (E.2) spells out the condition for the case of an exponential cost function. The left-hand side of equation (E.2) needs to be positive for $e$ to be positive. The condition for effort to be decreasing in $p$ is that $\ln(\pi(p+b))$ is bounded from above by $b/(p+b)$, which can be interpreted as the percent of the subjective belief in success due to overconfidence.  [^4] In the case of an exponential cost function, it is not necessary for effort to be low to generate a negative effect of $p$ on $W$. If $\gamma$ is low, e can be large even if $\ln(\pi(p+b))$ is only slightly positive.
E.2 The welfare effects of changes in $p$ and $b$
We model exposure to LLM input as a simulatneous change in $p$ and $b$. We now develop the Taylor approximation to $W$ to characterize the approximate effect of discrete changes in $p$ and $b$ on welfare.
As is common in these approximations, we assume that $e^{\prime\prime}(p+b)\approx 0$ and that these terms can therefore be neglected. We then calcuate the second-order approximation from the remaining terms to obtain
$W(p+\Delta p+b+\Delta b)-W(p,b)$ $\approx\ \pi\big{(}\underbrace{\left(e-be^{\prime}\right)\Delta p+\frac{e^{\prime}}{2}\cdot(\Delta p)^{2}}_{\text{ambiguous}}$ (8)
$\underbrace{-b\cdot e^{\prime}\Delta b+\frac{-e^{\prime}}{2}\cdot(\Delta b)^{2}}_{\text{negative}}\big{)}$
In order to evaluate a particular change in, both, $p$ and $b$, we need to make assumptions about their relative magnitude. Motivated in part by the empirical evidence from our experiment, we assume that $\Delta p=\Delta b>0$ due to exposure to LLM input.
Result 3. Assume that $\Delta b=\alpha\Delta p$ with $\alpha\geq 1$. The combined change reduces welfare if:
$\pi e<(1+\alpha)b\pi e^{\prime}$
In the case of an exponential cost function, this is the case if:
$\ln(\pi(p+b))<(1+\alpha)\frac{b}{p+b}$
The intuition behind Result 3 is most easily see for the case where $\Delta p$ and $\Delta b$ are approximately the same. The conditions for LLM exposure to increase welfare become even stricter than in Result 2. The reason is that, in addition to the dampening effect on the welfare gains from $p$, now the direct negative effects keep piling on. The case where the two effects are of the same magnitude is particularly simple, because the the second-order effects cancel.
If, more generally, $\Delta b=\alpha\Delta p$, where $\alpha>1$, the inequality would be even easier to satisfy, because the negative second-order effects from $\Delta b$ outweigh the positive ones from $\Delta p$, as can be seen in equation (8).
The final result provides a characterization of the welfare loss due to bias, i.e. it uses the Taylor approximation to calculate by how much welfare increased if $b=0$. This is not explicitly discussed in the text, but we report the quantity here for completeness.
Result 4: The welfare gains if bias were removed are given by
$\Delta W^{*}=\pi\frac{b^{2}e^{\prime}}{2}$
E.3 Piecing together the second derivatives for the Taylor approximation
As is customary in derive these approximations , we assume that behavior effects beyond the first order can be ignored, i.e. that $e^{\prime\prime}(p+b)=0$. Using this, we obtain
$\frac{\partial^{2}W}{\partial b^{2}}$ $=-\pi e^{2}$
$\frac{\partial^{2}W}{\partial p^{2}}$ $=\pi e^{2}$
For the cross-derivative, we obtain
$\frac{\partial}{\partial b}\left(\frac{\partial W}{\partial p}\right)=\frac{\partial^{2}W}{\partial b\partial p}=\pi\frac{\partial e}{\partial b}-\pi\frac{\partial e}{\partial b}=0.$
With the first derivatives defined earlier, we can then piece together the second-order Taylor approximation
$W(p+\Delta p+b+\Delta b)$ $\approx W(p+b)+\frac{\partial W}{\partial p}\cdot\Delta p+\frac{\partial^{2}W}{(\partial p)^{2}}\cdot\frac{1}{2}\cdot(\Delta p)^{2}$
$+\frac{\partial W}{\partial b}\cdot\Delta b+\frac{\partial^{2}W}{(\partial b)^{2}}\cdot\frac{1}{2}(\Delta b)^{2}$
$=W(p+b)+\left(\pi e-b\pi e^{\prime}\right)\Delta p+\frac{\pi e^{\prime}}{2}\cdot(\Delta p)^{2}$
$-b\pi\cdot e^{\prime}\Delta b+\frac{-\pi e^{\prime}}{2}\cdot(\Delta b)^{2}$